OpenStudy (anonymous):

integral(x+(x)^1/2)/(x)^2

OpenStudy (amistre64):

just split the fraction and run the powers

OpenStudy (anonymous):

\[\int\limits_{?}^{?} (x+\sqrt{x})\div X ^{2}\]

OpenStudy (anonymous):

so would the 1/x^2 be lnx

OpenStudy (amistre64):

x/x^2 = 1/x , which ints up to lnx yes what about the other one?

OpenStudy (amistre64):

x^(1/2 - 2) x^(1/2 - 2 + 1) -------------- (1/2 - 2 + 1)

OpenStudy (anonymous):

2/3x^3/2

OpenStudy (amistre64):

might be a negative in that someplace

OpenStudy (amistre64):

-.5 is what i see it as

OpenStudy (anonymous):

for my final im getting ln(x) - 2/(x)^1/2

OpenStudy (anonymous):

is this right

OpenStudy (amistre64):

+C, but yeah, i see it as that

OpenStudy (amistre64):

\[\int\frac{x+x^{1/2}}{x^2}dx\] \[\int\frac{x}{x^2}+\frac{x^{1/2}}{x^2}dx\] \[\int{x^{-1}}+x^{-3/2}dx=\ln|x|-2x^{-1/2}+C\]

OpenStudy (anonymous):

I just took a test so I am seeing how I did by posting questions here

OpenStudy (amistre64):

lol, ive done that in physics before

OpenStudy (anonymous):

heres one more integral 1/x(ln(x))^2

OpenStudy (anonymous):

its a good feeling when you see that your getting the correct answers

OpenStudy (amistre64):

\[\int \frac{1/x}{ln^2(x)}dx\] the top is practically the derivative of the bottom right? let u = ln^2(x) ; du = 2ln(x)* 1/x dx might be a suitable substitution if you go that route

OpenStudy (amistre64):

\[ln(ln^n(x))\to\frac{n}{x~ln(x)}\] sooo close

OpenStudy (amistre64):

-1/ln(x)

OpenStudy (amistre64):

\[-ln^{-1}(x)\] \[-(-1)ln^{-2}(x)~1/x=\frac{1}{x~ln^2(x)}\]

OpenStudy (amistre64):

i gotta go ... good luck

OpenStudy (anonymous):

thanks